OpenStudy (savannah_noelle):

I'm trying to figure out what the answer would be to this convex equation, can someone help me please? The convex lens has a focal length of 10.0cm, and the object distance is 15.0cm. The formula is 1/d(o) + 1/d(i) = 1/f. I got 0.06666667, is that correct?

OpenStudy (savannah_noelle):

@texaschic101 @Loser66 @hba can you guy help me?

OpenStudy (loser66):

I am sorry, I don't know :(

OpenStudy (savannah_noelle):

Awwuh okay :( well thank you for trying!

OpenStudy (jmcs1985):

You're halfway there....all you need to do is inverse (Good job, by the way!). \[\frac{ 1 }{ d(o) } = \frac{ 1 }{ 15 }, and \frac{1}{f} = \frac{1}{15}\] So, you now must input these values into the given equation: \[\frac{ 1 }{ 15 }+\frac{ 1 }{ d(i)} = \frac{ 1 }{10}\] Now solve algebraically: \[\frac{ 1 }{ d(i) }=\frac{1}{10}-\frac{1}{15} \rightarrow \frac{15}{150}-\frac{10}{150}=\frac{5}{150}=\frac{1}{30}\] Remember that your answer is inversed, so \[\frac{1}{30}\] is actually 30cm

OpenStudy (jmcs1985):

MCAT question for all fellow premed people like me ;)

OpenStudy (jmcs1985):

If you want to compare your answer to the one I provided, just divide 1 by 30, and you'll get 0.0666....

OpenStudy (savannah_noelle):

thank you so much! <3 I have one more question. How do you know the characteristics of a convex and concave lens? Also, do concave and convex lenses use the same equations?

OpenStudy (jmcs1985):

Excellent questions, @Savannah_Noelle. You're really giving me a run for the medals! 1. Concave lenses and convex lenses differ in the following fashion (see below): *Concave/converging lenses are thicker at the center and help magnify objects, Convex/diverging lenses are thinner at the center and also help (slightly) to magnify light but actually refract (bend) the light and not focus it at one point. 2. Convex and and concave lenses use the same equation as given above. Hope this helps! Feel free to let me know if you need more clarification. JMC

OpenStudy (jmcs1985):